A 0.01kg bullet is shot upward (+y axis) at 3600kph and hits and passes through the center of mass of a 5kg stationary block. The bullet emerges from the block and continues to move upward at 1440kph. To what maximum height does the block then rise above from its original position?
1
Expert's answer
2020-11-12T09:34:44-0500
"v^2=2gh\\\frac{1440}{3.6}=\sqrt{2(9.8)h}\\h=8163\ m"